Let the probability of occurrence of first event A, be ' a ' i..e., P(A)=a ∴P( not A)=1−a And also suppose that probability of occurrence of second event B,P(B)=b, ∴P(notB)=1−b Now, P(A and not B)+P( not A and B)=4926 ⇒P(A)×P(notB)+P(notA)×P(B)=4926⇒a×(1−b)+(1−a)b=4926⇒a+b−2ab=4926 And P( not A and not B)=4915 ⇒P(notA)×P( not B)=4915⇒(1−a)×(1−b)=4915⇒1−b−a+ab=4915⇒a+b−ab=4934 From (i) and (ii), a+b=4942 and ab=498 (a−b)2=(a+b)2−4ab=4942×4942−494×8=2401196 ∴a−b=4914 From (iii) and (iv), a=74,b=72 Hence probability of more probable of the two events =74
Given two independent events, if the probability that exactly one of them occurs is 26 49 and the probability that none of them occurs is 15 49, then…
JEE Main 2013 — Mathematics Probability & Statistics
Given two independent events, if the probability that exactly one of them occurs is 4926 and the probability that none of them occurs is 4915, then the probability of more probable of the two events is :
Held on 22 Apr 2013 · Verified 6 Jul 2026.
Options
- A
4/7
- B
6/7
- C
3/7
- D
5/7
Did you get this right?
Sign in to track your attempts and accuracy.
Your note
Sign in to keep a private note on this question. Nothing you write is ever public.
JEE Main Probability & Statistics in other years
More JEE Main Probability & Statistics questions
If the mean of the data <table class="pyq-table"><tbody><tr><th>Class</th><th>$5-10$</th><th>$10-15$</th><th>$15-20$</th><th>$20-25$</th><th>$25-30$</th><th>$30-35$</th></tr><tr><td>Frequency</td><td>$2$</td><td>$k$</td><td>$28$</td><td>$54$</td><td>$k+1$</td><td>$5$</td></tr></tbody></table> is $21$, then $k$ is one of the roots of the equation :
If the mean and the variance of the data \(\begin{array}{|c|c|c|c|c|} \hline \text{Class} & 4\text{-}8 & 8\text{-}12 & 12\text{-}16 & 16\text{-}20 \\ \hline \text{Frequency} & 3 & \lambda & 4 & 7 \\ \hline \end{array}\) are $\mu$ and 19 respectively, then the value of $\lambda+\mu$ is
The probability distribution of a random variable $X$ is given below : \(\begin{array}{|c|c|c|c|c|c|c|c|c|} \hline X & 4k & \dfrac{30}{7}k & \dfrac{32}{7}k & \dfrac{34}{7}k & \dfrac{36}{7}k & \dfrac{38}{7}k & \dfrac{40}{7}k & 6k \\ \hline P(X) & \dfrac{2}{15} & \dfrac{1}{15} & \dfrac{2}{15} & \dfrac{1}{5} & \dfrac{1}{15} & \dfrac{2}{15} & \dfrac{1}{5} & \dfrac{1}{15} \\ \hline \end{array}\) If $\mathrm{E}(\mathrm{X})=\frac{263}{15}$, then $\mathrm{P}(\mathrm{X}<20)$ is equal to :
A candidate has to go to the examination centre to appear in an examination. The candidate uses only one means of transportation for the entire distance out of bus, scooter and car. The probabilities of the candidate going by bus, scooter and car, respectively, are $\dfrac{2}{5}$, $\dfrac{1}{5}$ and $\dfrac{2}{5}$. The probabilities that the candidate reaches late at the examination centre are $\dfrac{1}{5}$, $\dfrac{1}{3}$ and $\dfrac{1}{4}$ if the candidate uses bus, scooter and car, respectively. Given that the candidate reached late at the examination centre, the probability that the candidate travelled by bus is:
A bag contains 6 blue and 6 green balls. Pairs of balls are drawn without replacement until the bag is empty. The probability that each drawn pair consists of one blue and one green ball is :
Other Mathematics topics for JEE Main
Keep practicing Probability & Statistics
Work through every JEE Main Probability & Statistics PYQ, year by year.